OpenStudy (jazzyluvsu123):

theres a number if you add 5 to it then divide that total by 3 you get 12. Whats the number? MEDAL AND FAN help plz

OpenStudy (anonymous):

31

OpenStudy (rockstar0765):

DO NOT GIVE DIRECT ANSWERS FOR THE LAST TIME EVERYONE

OpenStudy (jazzyluvsu123):

.-. Plz explain

OpenStudy (rockstar0765):

@JazzyLuvsU123 what do you think it is?

OpenStudy (anonymous):

do the math backwards. multiply 12 by three and subtract 5... It's simple after you learn to do it backwards

OpenStudy (fortytherapper):

Or set up an equation: Y = Total While you're looking for X 12 = x+5 ---- 3 Since you wanted to add 5 to that x then divide that number you find by 3 And Rockstar, you can be the police all you want, but people will continue to give the answers directly

OpenStudy (jazzyluvsu123):

@FortyTheRapper i still don't understand..

OpenStudy (jazzyluvsu123):

Like i set up the equation but I'm still not getting it

OpenStudy (fortytherapper):

You would need to get X by itself by doing the opposite thing to both sides. Since you are dividing by 3, what would you need to do to both sides to get rid of the division?

OpenStudy (jazzyluvsu123):

x+5/3= 12/3 ?

OpenStudy (vijeya3):

hmmmm...let the number be x if u are asked to add 5 to that unknown number,what would be the result? It wud be (x+5) Then your number is (x+5) Now you are asked to divide this number by 3 so now it becomes..... (x+5)/3 This newest number=12 that is it's value is 12 So....(x+5)/3=12 (x+5)=12*3 x+5=36 x=36-5 x=31 Hope this is helpful If it is..fan and medal me:)

OpenStudy (jazzyluvsu123):

Thanks SO much

OpenStudy (vijeya3):

never mind...u can always tag me when u have a math problem:)